OpenStudy (anonymous):

Can anyone answer... Find the equation of the tangent line to the graph of f at the point (-2,8)

OpenStudy (anonymous):

What is the graph f?

OpenStudy (anonymous):

y-8=f'(-2)*(x+2)

OpenStudy (anonymous):

What is the equation which describes f(x) ?

OpenStudy (anonymous):

the equation of the tangent of the graph of the function f(x) at (-2,8) [it must be 8=f(-2) ]is : \[y-8=f'(x)(x+2)\] such that \[f'(0)=[f(0)-8]/2\]

OpenStudy (zepp):

To do this question, you'll need to take the derivative of the function f(x), which is the tangent line.

OpenStudy (zepp):

Although you'd have to give us the equation that describes \(f(x)\)